# Design Stability in Adaptive Experiments: Implications for Treatment Effect Estimation

### v1

@@ 5.2 Efron’s Biased Coin Design
Moving beyond Wei’s adaptive coin design, we consider the biased coin design introduced by [3] which enforces another form of approximate balance between the number of allocations in the treatment and control groups. As in the previous section, let $D_{k}=m_{k}-n_{k}$ denote the imbalance between the treatment and control groups after the assignment of the $k$th unit, where $m_{k}$ and $n_{k}$ denote the numbers of treatment and control assignments, respectively. Under Efron’s biased coin design, the $i$th unit is assigned to treatment with probability:
@@
where $\eta\in[\frac{1}{2},1)$ controls the strength of the bias toward balance. In words, a larger value of $\eta$ forces faster correction of imbalance.
It is worth noting that $p_{i}$ takes the values $\eta$ and $1-\eta$ infinitely often, and thus Efron’s biased coin design is not strongly stable in the sense of Definition 2. However, the following lemma establishes weak stability of the design, making Theorems 1, 3, 4, and 6 applicable.
@@
Efron’s biased coin design (36) is weakly stable in sense of Definition 3, with $p_{1}^{\star}=\frac{4\eta^{2}(1-\eta)}{1-4\eta+12\eta^{2}-8\eta^{3}}$ and $p_{2}^{\star}=\frac{1-4\eta+8\eta^{2}-4\eta^{3}}{1-4\eta+12\eta^{2}-8\eta^{3}}$. Moreover,
$\frac{1}{N}\sum_{i=1}^{N}p_{i}\xrightarrow{p}\frac{1}{2}.$
@@
The proof of Lemma 2 proceeds by studying the treatment–control imbalance sequence $\{D_{k}\}_{k\geq 1}$. We first establish that $\{D_{k}\}_{k\geq 1}$ forms an irreducible and positively recurrent Markov chain by applying Foster’s Theorem [39]. The resulting positive recurrence and irreducibility ensure the existence of a unique stationary distribution, which, together with the mean ergodic theorem, facilitates characterization of the limiting behavior of long-run averages of functions of the assignment probabilities. Details of the proof are provided in Supplementary material 8.2.
@@
Furthermore, as Lemma 2 shows, under this design $N^{-1}\sum_{i=1}^{N}p_{i}\xrightarrow{p}\tfrac{1}{2}$, satisfying the sufficient condition in (18). This result allows for consistent estimation of $m^{2}_{0}$ and $m^{2}_{1}$ from (19) and (20), and of $\sigma^{2}_{0}$ and $\sigma^{2}_{1}$ from (28) and (29). Substituting these estimates into (17) and (30) yields conservative estimators of $V^{\text{IPW}}_{\text{Efron}}$ (37a) and $V^{\text{AIPW}}_{\text{Efron}}$ (37b):
@@
(a) IPW estimator under Non-Additive case
(b) AIPW estimator under Non-Additive case
(c) IPW estimator under Additive case
(d) AIPW estimator under Additive case
(e) IPW estimator under Log-Additive case
(f) AIPW estimator under Log-Additive case

### v2

@@ 5.2 Efron’s Biased Coin Design
Next, we consider the biased coin design introduced by [Efron1971-yh] which enforces another form of approximate balance between the number of allocations in the treatment and control groups. As in the previous section, let $D_{k}=m_{k}-n_{k}=2m_{k}-k$ denote the imbalance between the treatment and control groups after the assignment of the $k$th unit, where $m_{k}$ and $n_{k}$ denote the numbers of treatment and control assignments, respectively. Under Efron’s biased coin design, the $i$th unit is assigned to treatment with probability:
@@
where $\eta\in\big[\frac{1}{2},1)$ controls the strength of the bias toward balance. In words, a larger value of $\eta$ forces faster correction of imbalance. It is worth noting that $p_{i}$ takes the values $\eta$ and $1-\eta$ infinitely often, and thus Efron’s biased coin design is not strongly stable in the sense of Definition 2. However, the following lemma establishes weak stability of the design, making Theorems 1, 3, 4, and 6 applicable.
@@
Efron’s biased coin design (36) is weakly stable in sense of Definition 3, with $p_{1}^{\star}=\frac{4\eta^{2}(1-\eta)}{1-4\eta+12\eta^{2}-8\eta^{3}}$ and $p_{2}^{\star}=\frac{1-4\eta+8\eta^{2}-4\eta^{3}}{1-4\eta+12\eta^{2}-8\eta^{3}}$. Moreover, $N^{-1}\sum_{i=1}^{N}p_{i}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\frac{1}{2}$.
@@
The proof of Lemma 2 proceeds by studying the treatment–control imbalance sequence $\{D_{k}\}_{k\geq 1}$. We first establish that $\{D_{k}\}_{k\geq 1}$ forms an irreducible and positively recurrent Markov chain by applying Foster’s Theorem [Foster1953-av]. The resulting positive recurrence and irreducibility ensure the existence of a unique stationary distribution, which, together with the mean ergodic theorem, facilitates characterization of the limiting behavior of long-run averages of functions of the assignment probabilities. Details of the proof are provided in Section S9.2 of the Supplementary Material.
@@
Furthermore, as shown in Lemma 2, under this design $N^{-1}\sum_{i=1}^{N}p_{i}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\tfrac{1}{2}$, thereby satisfying the sufficient condition in (18). This allows for consistent estimation of $m^{2}_{0}$, $m^{2}_{1}$, $\sigma^{2}_{0}$, and $\sigma^{2}_{1}$. Substituting these into (17) and (30) yields conservative estimators of $V^{\text{IPW}}_{\text{Efron}}$ (37a) and $V^{\text{AIPW}}_{\text{Efron}}$ (37b):
@@
A complete assessment of the coverage of the confidence intervals constructed using the IPW and AIPW estimators under Efron’s design is provided in Section S7 of the Supplementary Material. The simulations use the same data-generating procedures and parameter settings as in Section 5.1, except that treatment assignments now follow Efron’s biased coin design (36) rather than Wei’s design. The biased-coin parameter is fixed at $\eta=0.7$ in all simulations.
Remark 11.
Both set of simulations conducted in Sections 5.1 and 5.2 reveal a common interesting pattern: when the nominal confidence level is high, the variance estimator appears less conservative, with empirical coverage closer to nominal. This phenomenon can be explained from the asymptotic theory. See Section S11 of the Supplementary Material for theoretical justification.
5.3 Possible extensions to more general designs in adaptive clinical trial literature and assessment of stability
As discussed in [Adaptive_clinical1] and [Adaptive_clinical2], Wei’s and Efron’s designs can be considered as special cases of more general classes of adaptive designs used in clinical trial literature, in which the assignment probability $p_{i}$ is a known function of the form $g\left(m_{i-1}/(i-1),\widehat{p}^{\star}_{i-1}\right)$, where as earlier, $m_{k}$ denotes the number of treatment units among the first $k$ subjects, and $\widehat{p}_{i-1}^{\star}$ is an estimator of the unknown target proportion $p^{\star}$ after stage (i1)(i-1). For designs in which $g(x,y)$ only depends on $x$ (e.g., Wei’s design with g(x,y)=1xg(x,y)=1-x), it is always possible to simulate sequences of assignment probabilities $p_{i}$ and plot them against step $i$ to see whether stability is attained (refer to Section S12 of the Supplementary Material for additional details). For designs where the target proportion $p^{\star}$ is unknown, but a simple estimator of the form (16) is used, such a simulation-based assessment is also possible. However, when the true target proportion is a known function of the unknown potential outcomes (e.g., in Neymanian allocation as in [Ravi2024]) and the estimated proportion is a function of the observed outcomes, imposing additional conditions on the potential outcomes may be necessary to conduct such simulation studies. This requires further investigation and should be an interesting area for further exploration.
6 Discussion
We have developed a general theoretical framework for conducting inference on average treatment effects in settings where treatment assignment is sequentially adaptive within a finite-population. This framework unifies and extends existing results by accommodating a broad class of adaptive randomization schemes, where assignment probabilities may evolve over time based on past outcomes. Within this setup, we establish central limit theorems for both inverse probability weighted (IPW) and augmented IPW (AIPW) estimators under strong and weak design stability conditions. Although the limiting distributions feature explicit expressions for the asymptotic variances, the fundamental problem of causal inference - not being able to observe the two potential outcomes for each unit - leads to challenges in their estimation. We propose conservative variance estimators that are consistent under different forms of treatment effect homogeneity.
From a practical standpoint, our findings provide reassurance that adaptive treatment assignment mechanisms—increasingly popular in modern experimental and clinical trial settings—can be used within a finite population framework without imposing any model on the potential outcome. The research opens up several new research possibilities. Extending the framework to covariate-adaptive designs where assignments depend explicitly on pre-measured covariates [[, e.g.,]]FSM1979, morgan2012, would broaden the applicability of the theory. Adaptive treatment assignment mechanisms also provide a natural solution to finding optimal designs in a finite-population setting, e.g., [Ravi2024] and the results presented in this paper can provide an inferential framework for such adaptive designs.
Acknowledgement
We thank the two reviewers and the Associate Editor for insightful and helpful comments. The research of Tirthankar Dasgupta was partially funded by National Science Foundation Grant SES 2217522. The research of Koulik Khamaru was partially funded by National Science Foundation Grant DMS-2311304.
S7 Simulation results for Efron’s biased coin design [Efron1971-yh]
As mentioned earlier, the simulations use the same data-generating procedures and parameter settings as in Section 5.1 of the main paper, except that treatment assignments now follow Efron’s biased coin design [Efron1971-yh] rather than Wei’s design [Wei1978-gd]. The biased-coin parameter is fixed at $\eta=0.7$ in all simulations.
(a) IPW estimator: Non-Additive
(b) IPW estimator: Additive
(c) IPW estimator: Log-Additive
(d) AIPW estimator: Non-Additive
(e) AIPW estimator: Additive
(f) AIPW estimator: Log-Additive